In a group of 78 people; 24 drink tea and 31 drink coffee. Find the probability that a person picked from this group at random is either tea or coffee drinker?
Find answer in fractions form.
1
Expert's answer
2020-09-09T19:42:19-0400

The probability that a person does not drink tea is "1-24\/78=54\/78." The probability that a person does not drink coffee is "1-31\/78=47\/78." Hence, the probability that a person does not drink coffee or tea equals "54/78\cdot47/78=54\cdot47/78^2=3\cdot47/(26\cdot13)=141/338." The probability that a person is tea or coffee drinker equals "1-141\/338=197\/338."